OpenStudy (anonymous):

Looking for someone who can work this out with me! I'm not looking for an answer, just help!! Part 1: Create and provide the dimensions for two similar figures of your choosing. Part 2: What is the similarity ratio of these figures along with the ratio of their surface area and volume? Part 3: Show your work, either using the actual volumes or using the formula, that the volume ratio is true.

OpenStudy (anonymous):

So far I made 2 similar equilateral cubes. One of them with the dimensions 9, 9, 9 and the other one with dimensions 27, 27, 27.

OpenStudy (anonymous):

9, 9, 9 Volume = 729^3 27, 27, 27 Volume = 19, 683^3 9, 9, 9 Surface Area = 486^2 27, 27, 27 Surface Area= 4,374^2

OpenStudy (anonymous):

Now how do I put this information (if it is right) into ratio form?

OpenStudy (phi):

What is the similarity ratio of these figures? ratio can be written as a fraction. in this case the similarity ratio is \[ \frac{9}{27} = \frac{1}{3} \] or it could be written as 1:3 (read as "one is to 3") do the same for the areas and volumes. If you did everything correctly, the areas should be in ratio 1:9 (1 to 3^2) and volume is 1:27 (1 to 3^3)

OpenStudy (phi):

btw, I would not write the answers as 9, 9, 9 Volume = 729^3 (because 729^3 means 729*729*729) the volume is 9*9*9 = 729 (cubic units (example cubic inches if you started with 9 inches for each side) 9*9*9= 9^3 (9^3 is just a short-cut for 9*9*9 ) and 9^3 = 729

OpenStudy (anonymous):

Thank you so much. :D I almost put that too.